OpenStudy (katielong):

how do you do reciprocals of decimals, e.g. 0.125?

OpenStudy (anonymous):

\[0.125=\frac{125}{1000} \]

OpenStudy (katielong):

how about 2.5 or 1.25?

OpenStudy (anonymous):

Refer to the attachment.

OpenStudy (agent0smith):

You can find the reciprocal of any number by dividing it into 1, such as... For any number n (whether it's a fraction, or decimal, or w/e) The reciprocal of n is \[\large \frac{ 1 }{ n }\] So if n = 0.125, then 1/n = 1/0.125 If n = 2.5, then 1/n = 1/2.5 If n = 1.25, then 1/n = 1/1.25

OpenStudy (katielong):

ah i see, thanks